• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name ipom

    Log entry time 00:00:35 on March 7,2007

    Entry number 195008

    keyword=shift summary

    16:00 Shift starts with no beam- waitng for MCC to return it, ~45 min
    16:47 Beam is back. taking a HARP scan at 5uA. going to empty target, raster OFF, spot size result: 131um 202um, repeated the scan and called MCC, they are working on it.
    17:28 MCC called to say that we might have did the HARP scan with a raster on even thogh the oscilloscpe display shows the raster is off.
    So they will run a HARP scan themself now.
    17:49 MCC agrees that they have a problem with the vertical beam width and it will take them a few minutes to fix it.
    19:15 MCC notified us that we got HV off on ch 0 and 1 on the beamline. we restated them. Jay from MCC is trying to reach the RC to discuss the vertical beam width problem that they cant seem to fix.
    21:28 MCC is saying that they managed to give a V:110um H:220um beam. and will now run tests on a higher current, we will consult with the RC
    21:38 run a HARP scan, at 5uA H:221 v:148
    23:09 took another HARP scan, the scan itself goes normaly, but when we start the analysis we get the message "Scan code died.check for bad data". RC said to ask MCC to take care of that. In the mean time MCC posted the result of their HARP scan at the Elog, entry No.1362724 which is H:236um V:109um. RC says we can go with that,but MCC has a few more tests they'd like to run.

    23:32 HARP scan done by MCC shows H:60um V:91um (elog entery 1362734)
    we'll go to production, going back to Pb.

    1218 - HARP scan 131um 202um
    1219 - HARP scan 131um 202um
    1220 - Junk
    1221 - HARP scan
    1222 - HARP scan
    1223 - HARP scan
    1224 - HARP scan
    1225 - Production Pb 53uA(max) Raster on. ps1=20 ps3=10 ps5=1